# From Pixels to Prose: Advancing Multi-Modal Language Models for Remote Sensing

### v1

@@ III-A1 Scene Description
Recent research has focused on improving the semantic understanding and caption quality of remote sensing images. Li et al. (2024) proposed a two-stage approach that combines cross-modal retrieval with semantic refinement [47]. This method addresses limitations in semantic understanding by extracting supplementary information from related remote sensing tasks, resulting in more accurate and contextually relevant image descriptions.
The integration of large language models (LLMs) with remote sensing image captioning has shown promising results. Silva et al. (2024) explored the use of LLMs for both image captioning and text-image retrieval tasks [34]. Their approach incorporates semantic segmentation techniques to enhance the model’s ability to identify and describe distinct regions within satellite imagery. This integration allows for a more comprehensive understanding of the scene, capturing both visual and semantic information.
To further improve captioning accuracy, researchers have developed consensus-aware semantic knowledge approaches. Li et al. (2024) introduced a method that utilizes multi-modal representations based on semantic-level relational features and visual-semantic contextual vectors [48]. This approach enables the model to generate more coherent and contextually appropriate captions by considering the consensus among different semantic aspects of the image.
Structured attention mechanisms have also been applied to high-resolution remote sensing image captioning. Zhao et al. (2021) developed a model that incorporates semantic segmentation techniques within a structured attention framework [49]. This approach allows for more precise identification of distinct regions and their characteristics, leading to improved captioning performance for complex, high-resolution satellite imagery.
@@ III-A2 Object Detection
Object detection and instance segmentation play a pivotal role in remote sensing for urban monitoring and environmental assessment by enabling the identification and delineation of entities within satellite and aerial imagery, including buildings, vehicles, and natural features. Techniques such as U-net-id have been effective in extracting detailed building information for urban planning and monitoring, aiding in assessing land use and infrastructure development [40]. Additionally, segmenting multi-temporal high-resolution imagery facilitates object-based change detection for tracking urban expansion and unauthorized construction [41]. Deep CNNs address challenges in detecting smaller objects while maintaining high accuracy, with advancements allowing pixel-scale uncertainty quantification [42]. Further enhancements from models like ISDNet contribute to detailed scene analysis for applications in traffic monitoring and green space assessment [43].
Recent advancements in MLLMs have further improved object detection. For instance, ContextDET employs a generate-then-detect framework that combines a visual encoder, a pre-trained language model, and a visual decoder to identify and locate objects within human vocabulary, demonstrating improved performance in contextual object detection tasks [44]. Similarly, VisionLLM utilizes a language-guided image tokenizer and an LLM-based decoder to seamlessly integrate various vision-centric tasks, including object detection and instance segmentation, into a unified model [45]. Additionally, LLM-Optic leverages large language models to enhance visual grounding by interpreting complex text queries and accurately identifying corresponding objects in images, achieving state-of-the-art zero-shot visual grounding capabilities [46]. These developments underscore the potential of MLLMs to advance object detection by effectively combining visual and textual information.
@@ III-A3 Change Detection
MLLMs enhance change detection in remote sensing by integrating visual and textual data. ChangeCLIP uses vision-language representation learning to identify surface changes from bitemporal images, achieving state-of-the-art performance on multiple datasets [37]. Similarly, LHRS-Bot utilizes volunteered geographic information and remote sensing images to perform various tasks, including change detection, demonstrating the potential of MLLMs in this domain [38]. Additionally, the development of a multimodal framework for remote sensing image change retrieval and captioning has enabled continuous monitoring of landscape changes, facilitating environmental protection and disaster monitoring [39]. These advancements underscore the efficacy of MLLMs in analyzing temporal sequences of satellite imagery to identify and characterize surface modifications over time.
@@ III-B1 Text-to-Image Retrieval
Text-to-image retrieval in remote sensing enables users to search satellite imagery databases using natural language queries. This cross-modal search aligns textual descriptions with visual features, facilitating efficient access to relevant Earth observation data. Recent advancements have focused on developing models that bridge the semantic gap between text and images. For instance, Yuan et al. proposed a method that integrates global and local information to enhance retrieval accuracy, addressing the challenge of capturing both overall scene context and specific object details [36].
To improve retrieval performance, some approaches incorporate external knowledge. Mi et al. introduced a knowledge-aware text-image retrieval method that enriches textual queries with information from external knowledge graphs, thereby reducing information asymmetry between modalities [4]. Additionally, Pan et al. presented a prior instruction representation learning paradigm, leveraging prior knowledge to instruct adaptive learning of vision and text representations, which enhances retrieval effectiveness [10].
Efforts have also been made to optimize retrieval efficiency. Mikriukov et al. developed a deep unsupervised contrastive hashing method for large-scale cross-modal text-image retrieval, enabling fast and memory-efficient searches without requiring labeled training samples [11]. These advancements collectively contribute to more accurate and efficient text-to-image retrieval systems in remote sensing, facilitating better utilization of satellite imagery for various applications.
@@ III-B2 Image-to-Text Generation
Automated caption generation for satellite imagery translates visual data into descriptive text, enhancing the accessibility and interpretability of remote sensing information. Recent advancements have employed transformer-based architectures to improve the quality of generated captions. Shen et al. introduced a model that integrates transformers with reinforcement learning, effectively capturing semantic information and addressing overfitting issues associated with limited datasets [32].
Incorporating attention mechanisms has further refined image-to-text generation. Li et al. proposed a multi-level attention model that mimics human cognitive processes by focusing on different image regions and corresponding textual elements, resulting in more accurate and contextually relevant descriptions [33]. Additionally, Silva et al. explored the application of large language models for captioning and retrieving remote sensing images, demonstrating that pre-trained models can be adapted to generate detailed and coherent captions for satellite imagery [34].
Visual grounding techniques play a crucial role in ensuring that generated captions accurately reflect the content and spatial relationships within imagery. Zhan et al. introduced the task of visual grounding for remote sensing data, aiming to localize referred objects in images based on natural language descriptions. Their work includes the development of a large-scale benchmark dataset and a transformer-based model that leverages multi-level cross-modal feature learning to enhance grounding accuracy [35]. These advancements collectively contribute to more precise and informative caption generation systems for satellite imagery.
@@ III-B3 Visual Question Answering
Visual Question Answering (VQA) systems in remote sensing enable users to query satellite imagery using natural language, facilitating interactive analysis of Earth observation data. These systems integrate visual and textual information to provide contextually relevant responses. Lobry et al. introduced the RSVQA dataset, which pairs remote sensing images with corresponding questions and answers, serving as a benchmark for developing and evaluating VQA models in this domain [28].
To enhance the performance of VQA systems, researchers have explored various model architectures. Zheng et al. proposed the Mutual Attention Inception Network (MAIN), which employs mutual attention mechanisms to effectively capture interactions between image features and textual queries, improving answer accuracy [29]. Similarly, Siebert et al. developed a Multi-Modal Fusion Transformer that leverages transformer-based architectures to fuse visual and textual modalities, demonstrating improved performance on VQA tasks [30].
Recent advancements have focused on lightweight models to reduce computational requirements. Hackel et al. introduced LiT-4-RSVQA, a transformer-based architecture designed for efficient VQA in remote sensing. This model achieves competitive accuracy while significantly reducing computational demands, making it suitable for operational scenarios [31]. These developments collectively contribute to more effective and accessible VQA systems for remote sensing applications.
@@ IV-A Benchmark Datasets
Benchmark datasets are essential for developing and evaluating MLLMs in remote sensing. The RSICap dataset, introduced by Hu et al., comprises 2,585 high-resolution remote sensing images, each annotated with five detailed captions. This dataset facilitates tasks such as image captioning and visual question answering by providing rich semantic information [27].
To address the need for larger datasets, Zhang et al. developed RS5M, a collection of 5 million remote sensing images paired with English descriptions. This dataset was created by filtering and captioning existing datasets, enabling the fine-tuning of vision-language models for tasks like zero-shot classification and cross-modal retrieval [9].
Another significant contribution is ChatEarthNet, presented by Yuan et al., which offers 163,488 image-text pairs with captions generated by ChatGPT-3.5, along with an additional 10,000 pairs from ChatGPT-4V. This dataset provides high-quality natural language descriptions for global-scale satellite data, enhancing the training and evaluation of vision-language models in remote sensing [26]
@@ IV-B Training Resources
The development of remote sensing MLLMs has been significantly advanced by the availability of pre-trained models and specialized training frameworks. These resources reduce computational demands and expedite the creation of new applications. SkySense is a multi-modal remote sensing foundation model pre-trained on a large-scale dataset comprising 21.5 million temporal sequences from high-spatial-resolution optical images, medium-resolution temporal multispectral imagery, and temporal synthetic aperture radar imagery [3]. SkySense’s architecture includes a factorized multi-modal spatiotemporal encoder, enabling it to handle diverse remote sensing tasks effectively.
In addition to pre-trained models, specialized training frameworks have been developed to support multi-modal learning in remote sensing. The Language-Guided Visual Prompt Compensation framework addresses modality absence by integrating language guidance into visual prompt learning, enhancing the model’s robustness in scenarios where certain modalities are missing [24]. This approach utilizes the complementary nature of language and visual data, improving the adaptability of multi-modal systems.
Furthermore, self-supervised learning techniques have been employed to pre-train models on multi-modal remote sensing data, reducing the reliance on labeled datasets. Berg et al. proposed a joint multi-modal self-supervised pre-training method that learns representations from unlabelled data, facilitating downstream tasks such as methane source classification [25]. This method demonstrates the potential of self-supervised learning in extracting meaningful features from multi-modal remote sensing data, thereby enhancing the efficiency and effectiveness of model training.
@@ IV-C Evaluation Metrics
Evaluating the performance of MLLMs, particularly in generating textual descriptions from satellite imagery, relies on several established metrics. The BLEU (Bilingual Evaluation Understudy) score assesses the precision of n-grams in the generated text against reference descriptions, providing insight into the model’s ability to produce accurate and fluent language [23]. Similarly, the METEOR (Metric for Evaluation of Translation with Explicit ORdering) metric evaluates both precision and recall, incorporating synonymy and stemming to better align with human judgment [22].
The CIDEr (Consensus-based Image Description Evaluation) metric is tailored for image captioning tasks. It measures the similarity of a generated description to a set of reference captions by considering term frequency-inverse document frequency (TF-IDF) weighting, thereby emphasizing the importance of informative words [21]. This approach ensures that the generated text aligns closely with human descriptions, capturing the nuances of the visual content.
These metrics have been applied to assess the quality of descriptions generated from satellite images. For instance, the EM-VLM4AD model, designed for visual question answering in autonomous driving, was evaluated using BLEU-4, METEOR, and CIDEr scores, demonstrating the applicability of these metrics in multi-modal tasks [70]. Such evaluations are crucial for developing models that effectively bridge the gap between visual data and natural language, enhancing the interpretability and usability of satellite imagery.
@@ V-A1 Computational Requirements
Processing high-resolution satellite imagery with MLLMs requires substantial computational resources. The integration of diverse data modalities, such as text and images, increases the complexity of these models, leading to higher computational demands. For instance, the training of large-scale MLLMs like LLaVA and multi-modal Phi-2 necessitates significant GPU resources to handle the extensive data and model parameters involved [12].
Memory constraints further limit the deployment of MLLMs in practical applications. High-resolution satellite images contain vast amounts of data, and processing them alongside textual information requires models with large memory capacities. To address this issue, researchers have explored parameter-efficient fine-tuning (PEFT) techniques. Context-PEFT, for example, introduces a novel framework that reduces the number of trainable parameters and GPU memory consumption without compromising performance [13].
Efforts to optimize MLLMs for resource-constrained environments have also been made. EM-VLM4AD is an efficient, lightweight, multi-frame vision-language model designed for autonomous driving applications. This model requires significantly less memory and computational power while achieving higher performance metrics compared to previous approaches, demonstrating the potential for deploying MLLMs in resource-limited settings [14]. These advancements highlight the ongoing efforts to balance computational requirements and model performance in the development of MLLMs for remote sensing applications.
@@ V-A2 Model Scalability
Scaling MLLMs to handle increasing data volumes and higher resolutions presents significant technical challenges. The relationship between model size and performance is complex, influencing practical deployment decisions. Recent studies have explored this relationship, revealing that while larger models often achieve better performance, they also require more computational resources. For instance, Aghajanyan et al. investigated scaling laws for generative mixed-modal language models, demonstrating that performance improvements diminish as model size increases, highlighting the need for efficient scaling strategies [15].
To address these challenges, researchers have proposed various approaches to enhance model scalability. One such method is multi-resolution training, which allows models to process inputs at different resolutions, thereby improving efficiency without compromising performance. Tian et al. introduced ResFormer, a model that employs multi-resolution training to scale vision transformers effectively, enabling them to handle high-resolution inputs with reduced computational costs [16]. Additionally, integrating multi-modal data poses unique scalability issues, as models must effectively process and align information from diverse sources. Aghajanyan et al. also examined scaling laws for generative mixed-modal language models, providing insights into how model size and data complexity interact in multi-modal settings [15].
Efforts to improve scalability have also focused on optimizing model architectures and training strategies. For example, Liu et al. developed SPHINX-X, a family of multi-modal large language models that scale data and parameters efficiently. Their approach involves modifying the model architecture to remove redundant components and simplifying the training process, resulting in improved scalability and performance [17]. These advancements underscore the importance of developing scalable MLLMs capable of handling large-scale, high-resolution data in practical applications.

### v2

@@ III-A1 Scene Description
Recent research has focused on improving the semantic understanding and caption quality of remote sensing images. Li et al. (2024) proposed a two-stage approach that combines cross-modal retrieval with semantic refinement [li2024cross]. This method addresses limitations in semantic understanding by extracting supplementary information from related remote sensing tasks, resulting in more accurate and contextually relevant image descriptions.
The integration of large language models (LLMs) with remote sensing image captioning has shown promising results. Silva et al. (2024) explored the use of LLMs for both image captioning and text-image retrieval tasks [silva2024large]. Their approach incorporates semantic segmentation techniques to enhance the model’s ability to identify and describe distinct regions within satellite imagery. This integration allows for a more comprehensive understanding of the scene, capturing both visual and semantic information.
To further improve captioning accuracy, researchers have developed consensus-aware semantic knowledge approaches. Li et al. (2024) introduced a method that utilizes multi-modal representations based on semantic-level relational features and visual-semantic contextual vectors [li2024learning]. This approach enables the model to generate more coherent and contextually appropriate captions by considering the consensus among different semantic aspects of the image.
Structured attention mechanisms have also been applied to high-resolution remote sensing image captioning. Zhao et al. (2021) developed a model that incorporates semantic segmentation techniques within a structured attention framework [zhao2021high]. This approach allows for more precise identification of distinct regions and their characteristics, leading to improved captioning performance for complex, high-resolution satellite imagery.
@@ III-A2 Object Detection
Object detection and instance segmentation play a pivotal role in remote sensing for urban monitoring and environmental assessment by enabling the identification and delineation of entities within satellite and aerial imagery, including buildings, vehicles, and natural features. Techniques such as U-net-id have been effective in extracting detailed building information for urban planning and monitoring, aiding in assessing land use and infrastructure development [wagner2020u]. Additionally, segmenting multi-temporal high-resolution imagery facilitates object-based change detection for tracking urban expansion and unauthorized construction [zhang2017separate]. Deep CNNs address challenges in detecting smaller objects while maintaining high accuracy, with advancements allowing pixel-scale uncertainty quantification [kampffmeyer2016semantic]. Further enhancements from models like ISDNet contribute to detailed scene analysis for applications in traffic monitoring and green space assessment [garg2021isdnet].
Recent advancements in MLLMs have further improved object detection. For instance, ContextDET employs a generate-then-detect framework that combines a visual encoder, a pre-trained language model, and a visual decoder to identify and locate objects within human vocabulary, demonstrating improved performance in contextual object detection tasks [zang2024contextual]. Similarly, VisionLLM utilizes a language-guided image tokenizer and an LLM-based decoder to seamlessly integrate various vision-centric tasks, including object detection and instance segmentation, into a unified model [wang2024visionllm]. Additionally, LLM-Optic leverages large language models to enhance visual grounding by interpreting complex text queries and accurately identifying corresponding objects in images, achieving state-of-the-art zero-shot visual grounding capabilities [zhao2024llm]. These developments underscore the potential of MLLMs to advance object detection by effectively combining visual and textual information.
@@ III-A3 Change Detection
MLLMs enhance change detection in remote sensing by integrating visual and textual data. ChangeCLIP uses vision-language representation learning to identify surface changes from bitemporal images, achieving state-of-the-art performance on multiple datasets [dong2024changeclip]. Similarly, LHRS-Bot utilizes volunteered geographic information and remote sensing images to perform various tasks, including change detection, demonstrating the potential of MLLMs in this domain [muhtar2024lhrsbot]. Additionally, the development of a multimodal framework for remote sensing image change retrieval and captioning has enabled continuous monitoring of landscape changes, facilitating environmental protection and disaster monitoring [ferrod2024towards]. These advancements underscore the efficacy of MLLMs in analyzing temporal sequences of satellite imagery to identify and characterize surface modifications over time.
@@ III-B1 Text-to-Image Retrieval
Text-to-image retrieval in remote sensing enables users to search satellite imagery databases using natural language queries. This cross-modal search aligns textual descriptions with visual features, facilitating efficient access to relevant Earth observation data. Recent advancements have focused on developing models that bridge the semantic gap between text and images. For instance, Yuan et al. proposed a method that integrates global and local information to enhance retrieval accuracy, addressing the challenge of capturing both overall scene context and specific object details [yuan2022remote].
To improve retrieval performance, some approaches incorporate external knowledge. Mi et al. introduced a knowledge-aware text-image retrieval method that enriches textual queries with information from external knowledge graphs, thereby reducing information asymmetry between modalities [mi2024knowledge]. Additionally, Pan et al. presented a prior instruction representation learning paradigm, leveraging prior knowledge to instruct adaptive learning of vision and text representations, which enhances retrieval effectiveness [pan2024pir].
Efforts have also been made to optimize retrieval efficiency. Mikriukov et al. developed a deep unsupervised contrastive hashing method for large-scale cross-modal text-image retrieval, enabling fast and memory-efficient searches without requiring labeled training samples [mikriukov2022deep]. These advancements collectively contribute to more accurate and efficient text-to-image retrieval systems in remote sensing, facilitating better utilization of satellite imagery for various applications.
@@ III-B2 Image-to-Text Generation
Automated caption generation for satellite imagery translates visual data into descriptive text, enhancing the accessibility and interpretability of remote sensing information. Recent advancements have employed transformer-based architectures to improve the quality of generated captions. Shen et al. introduced a model that integrates transformers with reinforcement learning, effectively capturing semantic information and addressing overfitting issues associated with limited datasets [shen2020remote].
Incorporating attention mechanisms has further refined image-to-text generation. Li et al. proposed a multi-level attention model that mimics human cognitive processes by focusing on different image regions and corresponding textual elements, resulting in more accurate and contextually relevant descriptions [li2020multi]. Additionally, Silva et al. explored the application of large language models for captioning and retrieving remote sensing images, demonstrating that pre-trained models can be adapted to generate detailed and coherent captions for satellite imagery [silva2024large].
Visual grounding techniques play a crucial role in ensuring that generated captions accurately reflect the content and spatial relationships within imagery. Zhan et al. introduced the task of visual grounding for remote sensing data, aiming to localize referred objects in images based on natural language descriptions. Their work includes the development of a large-scale benchmark dataset and a transformer-based model that leverages multi-level cross-modal feature learning to enhance grounding accuracy [zhan2023rsvg]. These advancements collectively contribute to more precise and informative caption generation systems for satellite imagery.
@@ III-B3 Visual Question Answering
Visual Question Answering (VQA) systems in remote sensing enable users to query satellite imagery using natural language, facilitating interactive analysis of Earth observation data. These systems integrate visual and textual information to provide contextually relevant responses. Lobry et al. introduced the RSVQA dataset, which pairs remote sensing images with corresponding questions and answers, serving as a benchmark for developing and evaluating VQA models in this domain [lobry2020rsvqa].
To enhance the performance of VQA systems, researchers have explored various model architectures. Zheng et al. proposed the Mutual Attention Inception Network (MAIN), which employs mutual attention mechanisms to effectively capture interactions between image features and textual queries, improving answer accuracy [zheng2021mutual]. Similarly, Siebert et al. developed a Multi-Modal Fusion Transformer that leverages transformer-based architectures to fuse visual and textual modalities, demonstrating improved performance on VQA tasks [siebert2022multi].
Recent advancements have focused on lightweight models to reduce computational requirements. Hackel et al. introduced LiT-4-RSVQA, a transformer-based architecture designed for efficient VQA in remote sensing. This model achieves competitive accuracy while significantly reducing computational demands, making it suitable for operational scenarios [hackel2023lit]. These developments collectively contribute to more effective and accessible VQA systems for remote sensing applications.
@@ IV-A Benchmark Datasets
Benchmark datasets are essential for developing and evaluating MLLMs in remote sensing. The RSICap dataset, introduced by Hu et al., comprises 2,585 high-resolution remote sensing images, each annotated with five detailed captions. This dataset facilitates tasks such as image captioning and visual question answering by providing rich semantic information [hu2023rsgpt].
To address the need for larger datasets, Zhang et al. developed RS5M, a collection of 5 million remote sensing images paired with English descriptions. This dataset was created by filtering and captioning existing datasets, enabling the fine-tuning of vision-language models for tasks like zero-shot classification and cross-modal retrieval [zhang2023rs5m].
Another significant contribution is ChatEarthNet, presented by Yuan et al., which offers 163,488 image-text pairs with captions generated by ChatGPT-3.5, along with an additional 10,000 pairs from ChatGPT-4V. This dataset provides high-quality natural language descriptions for global-scale satellite data, enhancing the training and evaluation of vision-language models in remote sensing [yuan2024chatearthnet]
@@ IV-B Training Resources
The development of remote sensing MLLMs has been significantly advanced by the availability of pre-trained models and specialized training frameworks. These resources reduce computational demands and expedite the creation of new applications. SkySense is a multi-modal remote sensing foundation model pre-trained on a large-scale dataset comprising 21.5 million temporal sequences from high-spatial-resolution optical images, medium-resolution temporal multispectral imagery, and temporal synthetic aperture radar imagery [guo2024skysense]. SkySense’s architecture includes a factorized multi-modal spatiotemporal encoder, enabling it to handle diverse remote sensing tasks effectively.
In addition to pre-trained models, specialized training frameworks have been developed to support multi-modal learning in remote sensing. The Language-Guided Visual Prompt Compensation framework addresses modality absence by integrating language guidance into visual prompt learning, enhancing the model’s robustness in scenarios where certain modalities are missing [huang2024language]. This approach utilizes the complementary nature of language and visual data, improving the adaptability of multi-modal systems.
Furthermore, self-supervised learning techniques have been employed to pre-train models on multi-modal remote sensing data, reducing the reliance on labeled datasets. Berg et al. proposed a joint multi-modal self-supervised pre-training method that learns representations from unlabelled data, facilitating downstream tasks such as methane source classification [berg2023joint]. This method demonstrates the potential of self-supervised learning in extracting meaningful features from multi-modal remote sensing data, thereby enhancing the efficiency and effectiveness of model training.
@@ IV-C Evaluation Metrics
Evaluating the performance of MLLMs, particularly in generating textual descriptions from satellite imagery, relies on several established metrics. The BLEU (Bilingual Evaluation Understudy) score assesses the precision of n-grams in the generated text against reference descriptions, providing insight into the model’s ability to produce accurate and fluent language [papineni2002bleu]. Similarly, the METEOR (Metric for Evaluation of Translation with Explicit ORdering) metric evaluates both precision and recall, incorporating synonymy and stemming to better align with human judgment [banerjee2005meteor].
The CIDEr (Consensus-based Image Description Evaluation) metric is tailored for image captioning tasks. It measures the similarity of a generated description to a set of reference captions by considering term frequency-inverse document frequency (TF-IDF) weighting, thereby emphasizing the importance of informative words [vedantam2015cider]. This approach ensures that the generated text aligns closely with human descriptions, capturing the nuances of the visual content.
These metrics have been applied to assess the quality of descriptions generated from satellite images. For instance, the EM-VLM4AD model, designed for visual question answering in autonomous driving, was evaluated using BLEU-4, METEOR, and CIDEr scores, demonstrating the applicability of these metrics in multi-modal tasks [gopalkrishnan2023em]. Such evaluations are crucial for developing models that effectively bridge the gap between visual data and natural language, enhancing the interpretability and usability of satellite imagery.
@@ V-A1 Computational Requirements
Processing high-resolution satellite imagery with MLLMs requires substantial computational resources. The integration of diverse data modalities, such as text and images, increases the complexity of these models, leading to higher computational demands. For instance, the training of large-scale MLLMs like LLaVA and multi-modal Phi-2 necessitates significant GPU resources to handle the extensive data and model parameters involved [basu2024understanding].
Memory constraints further limit the deployment of MLLMs in practical applications. High-resolution satellite images contain vast amounts of data, and processing them alongside textual information requires models with large memory capacities. To address this issue, researchers have explored parameter-efficient fine-tuning (PEFT) techniques. Context-PEFT, for example, introduces a novel framework that reduces the number of trainable parameters and GPU memory consumption without compromising performance [hadji2023context].
Efforts to optimize MLLMs for resource-constrained environments have also been made. EM-VLM4AD is an efficient, lightweight, multi-frame vision-language model designed for autonomous driving applications. This model requires significantly less memory and computational power while achieving higher performance metrics compared to previous approaches, demonstrating the potential for deploying MLLMs in resource-limited settings [gopalkrishnan2024multi]. These advancements highlight the ongoing efforts to balance computational requirements and model performance in the development of MLLMs for remote sensing applications.
@@ V-A2 Model Scalability
Scaling MLLMs to handle increasing data volumes and higher resolutions presents significant technical challenges. The relationship between model size and performance is complex, influencing practical deployment decisions. Recent studies have explored this relationship, revealing that while larger models often achieve better performance, they also require more computational resources. For instance, Aghajanyan et al. investigated scaling laws for generative mixed-modal language models, demonstrating that performance improvements diminish as model size increases, highlighting the need for efficient scaling strategies [aghajanyan2023scaling].
To address these challenges, researchers have proposed various approaches to enhance model scalability. One such method is multi-resolution training, which allows models to process inputs at different resolutions, thereby improving efficiency without compromising performance. Tian et al. introduced ResFormer, a model that employs multi-resolution training to scale vision transformers effectively, enabling them to handle high-resolution inputs with reduced computational costs [tian2022resformer]. Additionally, integrating multi-modal data poses unique scalability issues, as models must effectively process and align information from diverse sources. Aghajanyan et al. also examined scaling laws for generative mixed-modal language models, providing insights into how model size and data complexity interact in multi-modal settings [aghajanyan2023scaling].
Efforts to improve scalability have also focused on optimizing model architectures and training strategies. For example, Liu et al. developed SPHINX-X, a family of multi-modal large language models that scale data and parameters efficiently. Their approach involves modifying the model architecture to remove redundant components and simplifying the training process, resulting in improved scalability and performance [liu2024sphinx]. These advancements underscore the importance of developing scalable MLLMs capable of handling large-scale, high-resolution data in practical applications.